Use the following information to answer questions 6 and 7:

An analyst gathered the following information regarding Diago Investments:

  • FCFF at the end of 2011 = $1.1 million
  • Interest expense = $525,000
  • Fixed capital expenditure = $650,000
  • Working capital expenditure = $280,000
  • Depreciation expense = $395,000
  • Net borrowing = $480,000
  • Number of common shares outstanding = 600,000
  • Weighted average cost of capital = 14%
  • Risk free rate of return = 5%
  • Equity market risk premium = 7%
  • Beta of the company's stock = 1.1
  • Expected long-term growth rate in FCFE = 6%
  • Tax rate = 40%

Question: FCFE per share at the end of 2011 is closest to:

Select one:

a. $2.99

b. $2.11

c. $3.82

Question 7

Question text

The value per share of the company's stock at the end of 2011 is closest to:

Select one:

a. $33.36

b. $27.93

c. $31.47
Answer #1

6

/shares outstanding

FCFE per share= (1100000-525000*(1-0.4)+480000)/600000

=2.11

7

As per CAPM
expected return = risk-free rate + beta * (Market risk premium)
Expected return% = 5 + 1.1 * (7)
Expected return% = 12.7
Price = FCFE* (1 + growth rate )/(cost of equity - growth rate)
Price = 2.11 * (1+0.06) / (0.127 - 0.06)
Price = 33.36